OpenStudy (anonymous):

complete square root add what number to both sides to x^2-6x=-8

OpenStudy (amistre64):

what is a complete square?

OpenStudy (amistre64):

add (b/2)^2

OpenStudy (anonymous):

I did that I need a whole number I came up with 1 but it is not right

OpenStudy (anonymous):

-6/2 = -3 then you square that (-3)^2 = 9

OpenStudy (amistre64):

blext didnt get a 1 ;)

OpenStudy (anonymous):

b=-6/2^2 = 36/4

OpenStudy (amistre64):

yes; and 36/4 = 9

OpenStudy (anonymous):

where so you get 36?

OpenStudy (amistre64):

6^2 = 36 --------- = 9 2^2 = 4

OpenStudy (anonymous):

formula x^-6x=-8 b=(-6/2)^2

OpenStudy (anonymous):

36/4 =9 so I put 9 on each side of the equation?

OpenStudy (amistre64):

welll close; b = -6; the number you wanna add to both sides is: (b/2)^2

OpenStudy (amistre64):

yes; add 9 to both sides to complete the square

OpenStudy (anonymous):

okay that is not the example in my book

OpenStudy (amistre64):

there are many ways to get to the same answer; you gotta pick what works for you :)

OpenStudy (anonymous):

this is true I am taking this test for my last time on my math lab and want to get the correct answers.

OpenStudy (amistre64):

the real solution is: (bx/2ax)^2 but for the a=1 thats rather obiquitous. and the x's always cancel anyways.

OpenStudy (anonymous):

thank you